OpenStudy (anonymous):

Write z1 and z2 in polar form. (Express θ in radians.) z1 = [(2)^1/2] − [(2)^1/2] i, z2 = 1 − i

OpenStudy (anonymous):

z1=\[\sqrt{2}-\sqrt{2}i\]

OpenStudy (anonymous):

|z1| = sqrt(sqrt(2)^2 + sqrt(2)^2) = sqrt(2+2) = sqrt(4) = 2 arg(z1) = arctan(-sqrt(2)/sqrt(2)) = arctan(-1) = -pi/4 or 3pi/4 Re(z1) = +ve, Im(z1) = -ve, therefore -pi/2 < arg(z1) < 0 therefore = -pi/4

OpenStudy (anonymous):

|z2| = sqrt(1^2 + 1^2) = sqrt(2) arg(z2) = arctan(-1/1) = arctan(-1) and then follows the same logic as for z1 so the argument is also -pi/4

OpenStudy (anonymous):

so overall: z1 = 2 cis -pi/4 z2 = sqrt(2) cis -pi/4

OpenStudy (anonymous):

this stuff really clicks for me now. awesome responses. thanks.